Molar Mass Al = 26.98 grams/mole

Molar Mass HCl = 36.46 grams/mole

Molar Mass AlCl3 = 133.33 grams/mole

Molar Mass H2 = 2.02 grams/mole

 

How many grams of AlCl3 will be made if you fully react 93 grams of HCl?

Step 1: Convert grams of HCl to moles of HCl

Step 2: Convert moles of HCl to moles of AlCl3

Step 3: Convert moles of AlCl3 to grams of AlCl3

1
Expert's answer
2022-01-23T23:35:03-0500

2Al(s)+6HCl(aq)→2AlCl3(aq)+3H2(g)

Step1: moles of HCl= (93/36.46)=2.55 moles

Step 2: mole ratio= 6:2= 3:1

2.55÷ 3= 0.85moles

Step 3: grams of AlCl3= 0.85×133.33= 113.33g